OpenStudy (anonymous):

PLEASE HELP WILL FAN AND MEDAL. Find the volume of a sphere with radius 9 yd.

OpenStudy (anonymous):

@IrishBoy123 @arindameducationusc

OpenStudy (anonymous):

hey may I jump in to give my help? :)

OpenStudy (irishboy123):

go for it mate!

OpenStudy (anonymous):

Sure.

OpenStudy (anonymous):

I got V≈2334.67m³ is it right ?

OpenStudy (anonymous):

ok I really hate math but I love to solve these I like the challenge.... anyways :) you would have you radius = 9 *3.14 which is pi then divide. one you do that I will say what to divide by. :) btw that is how I was taught the easiest way by far

OpenStudy (anonymous):

Well for some reason that's not one of the answer choices.

OpenStudy (irishboy123):

\[V = \frac{4}{3} \pi r^3\]

OpenStudy (anonymous):

Here are the options: A) 3,829.7 yd^3 B) 3,052.1 yd^3 C) 381.5 yd3^3 D) 339.1 yd^3

OpenStudy (anonymous):

@adajiamcneal @ganeshie8

OpenStudy (anonymous):

@Christian_10_

OpenStudy (anonymous):

@calculusxy

OpenStudy (calculusxy):

\[V = \frac{ 4 }{ 3 }\pi r^3 = \frac{ 4 }{ 3 }(3.14)(9)^3 \approx 3044.4498yd^3\]

OpenStudy (calculusxy):

@Darion17

OpenStudy (calculusxy):

That was the answer that I got.

OpenStudy (anonymous):

Thank you much!

OpenStudy (calculusxy):

You're welcome :)